# Overcoming the Zero-Rate Hashing Bound with Holographic Quantum Error Correction

### v2

@@ II.2 Tensor-Network Decoding
Contracting a tensor network representing a holographic-code state with potentially thousands of qubits is computationally intractable; as such, we follow the simplification proposed in [36, 37]. In what follows, we briefly review this technique.
@@
As a side remark, the tensors of Equation 1 are not, strictly speaking, isometric. However, they facilitate efficient contraction, provided that the proper index contraction sequence is indicated; as such, the full decoding process which we will describe below exhibits a runtime complexity of $\mathcal{O}(n^{2.37})$ in the best case [70, 71]. More details on the formation of large tensor networks using the aforementioned scheme can be found in [36, 72].
Tensor-network decoder is a maximum-likelihood (ML) decoder, which is shown to be optimal given the error model. The ML decoder generally works by calculating the most-likely correction needed to return the system to the correct code state, given a syndrome. Here, we define a syndrome as the set $\mathbf{s}=[s_{1}\cdots s_{n-k}]$ of $\mathbb{Z}_{2}$ parity results from measuring each stabilizer. An error $\mathsf{E}$ is initialized as a probability vector corresponding to each individual physical qubit; this vector is then contracted with the holographic tensor network, permitting us to calculate the syndrome $\mathbf{s}=\mathsf{HE}$, where $\mathsf{H}$ is the parity-check matrix of the holographic code.
@@
where $\text{Prob}[\mathsf{E}_{\text{pure}}|SL]$ signifies the probability of a pure error acting on the coset of stabilizer $S\in\mathsf{S}$ with respect to logical operator $L\in\mathsf{L}$. In practice, the decoder must find $L$ such that $\mathbb{P}(\mathsf{L},\mathbf{s})$ is maximized, i.e. $\operatorname*{arg\,max}\big{[}\mathbb{P}(\mathsf{L},\mathbf{s})\big{]}$. As an additional condition, the summation over all logical cosets must be equivalent to unity; that is $\sum_{L\in\mathsf{L}}\sum_{S\in\mathsf{S}}\mathbb{P}(\mathsf{L},\mathbf{s})=1$. As before, we refer the reader to [36, 37, 72] for further details on calculating the logical success rate.
@@ III.1 Noise Model and Setup
The error model tested in this work is principally the code-capacity error model for Pauli noise. Such an error model takes on the form
@@
HaPPY $XZZXI\bar{I}$, $IXZZX\bar{I}$, $XIXZZ\bar{I}$, $ZXIXZ\bar{I}$ $ZZZZZ\bar{Z},XXXXX\bar{X}$
@@
Steane $XXIIIXX\bar{I}$, $IXXXIIX\bar{I}$, $IIIXXXX\bar{I}$, $ZZIIIZZ\bar{I}$, $IZZZIIZ\bar{I}$, $IIIZZZZ\bar{I}$ $ZZZZZZZ\bar{Z}$, $XXXXXXX\bar{X}$
$\llbracket 6,1,3\rrbracket$ $ZIZIII\bar{I}$, $XZYYXI\bar{I}$, $XXXXZI\bar{I}$, $IZZXIX\bar{I}$, $XYXYIZ\bar{I}$ $XZXZII\bar{X}$, $XYYXII\bar{Z}$
SCF $XXIXI\bar{I}$, $IIXXX\bar{I}$, $ZIZZI\bar{I}$, $IZIZZ\bar{I}$ $XIXII\bar{X}$, $IIZIZ\bar{Z}$
CD-Steane $XZZIIIX\bar{I}$, $XIZXZII\bar{I}$, $XIIIZZX\bar{I}$, $ZXXIIIZ\bar{I}$,$ZIXZXII\bar{I}$, $ZIIIXXZ\bar{I}$ $XZZXZZX\bar{X}$, $ZXXZXXZ\bar{Z}$
@@
Threshold calculations were performed using the tensor-network decoder from [36, 37]. These decoders and others will be made public in an upcoming software package for holographic quantum error correction codes [73]. However, for the purposes of understanding the current work, we focus on surveying biased-noise resilience for a swath of known holographic codes, and not on a systematic treatment of tensor-network decoding methodology [70].
@@
In our biased-noise threshold profiling, we tested 43 distinct Pauli biases and have displayed them on the ternary diagram shown in Figure 2. Each point of the triangular plot shown represents a pure Pauli error channel. Subsequently, points on the interior of the plot can be read by following the grid lines provided to the boundaries; for example, the point to the lower left of the central depolarizing noise marker can be interpreted as having relative biases $\bar{r}=(r_{X},~{}r_{Y},~{}r_{Z})$ of $(0.25,~{}0.25,~{}0.5)$. At each bias point 10,000 Monte Carlo simulations were performed per threshold curve data point, per layer. We display commensurate threshold curve results for some select data points in Figure 4 for the zero-rate HaPPY code.
@@ III.2 Zero-Rate Holographic Code Results
Figure 3: Ternary plots for all holographic codes investigated in this study under biased noise, in the code-capacity setting. (a)-(d) depicts the zero-rate HaPPY, Steane, $\llbracket 6,1,3\rrbracket$, and SCF codes; thresholds are color-coded from dark blue ($p_{th}=10\%$) to dark red ($p_{th}=50\%$). Additionally, we tested the tailored $\llbracket 7,1,3\rrbracket$ code from [38, 24], which achieved an identical threshold profile to that of the HaPPY code.
Figure 4: Threshold curves for depolarizing, pure $X$, $Y$, and $Z$ noise, as studied using the tensor-network decoder, for the zero-rate HaPPY code at up to $L=3$ layers of edge inflation.
@@
As a more detailed example, we plotted in Figure 4 the individual threshold curves obtained for the zero-rate HaPPY code for depolarizing noise, as well as for biased pure $X$, pure $Y$, and pure $Z$ noise. Around the fixed points in subfigures (b)-(d), the recovery rate promptly increases, which is due to several reasons: Firstly, the pure 1-Pauli noise capacity can be shown to be $50\%$. Secondly, as the tensor-network decoder is a maximum-likelihood decoder, after the $p=0.50$ mark, the decoder selects the most-probable error and returns a pure error (destabilizer) vector which is used to correct. Therefore, unlike a minimum weight decoder, the tensor network decoder is still able to provide the right correction with high probability even when $p>0.50$. More details can be found in [40, 72].
If we look beyond the pure 1-Pauli portion of the ternary plots, we note more subtle behavior of these codes: for example, it can be seen that modest threshold increases emerge for the Steane and $\llbracket 6,1,3\rrbracket$ codes as we move towards pure 2-Pauli noise, i.e., $XY$, $YZ$, and $XZ$ noise. These changes can be seen more clearly by examining the threshold behavior of the codes via the tuning of a bias parameter $\eta$. In previous work, this aim was accomplished by comparing various zero-rate codes against the hashing bound, a useful lower bound for quantifying code-capacity channel performance [25]. However, in contrast to the present work, all previous works have examined the pure 1-Pauli noise limit associated with the hashing bound, and have not delved into combinations of two types of Pauli errors at once, a condition known to occur in the square-lattice GKP code under a symmetric Gaussian random displacement noise channel [74, 75, 76, 77], as well as for error channels resulting from performing Pauli twirling on the amplitude damping noise channel [78]. As such, we tune the $\eta$ parameter within the range $0\leq\eta\leq+\infty$, extending previous work past the regime indicated by $\eta\geq 0.5$ [25, 20, 24].
@@
where $R$ represents an achievable rate $k/n$ for a random stabilizer code and $H(\bar{p})$ represents the Shannon entropy [1]
$H(\bar{p})=-\sum_{i\in\{I,X,Y,Z\}}p_{i}\log{p_{i}}~{}.$ ) . (6)
@@
1,3,10,30,100,300,1000,+) 1,3,10,30,100,300,1000,$\infty$)\ .1 , 3 , 10 , 30 , 100 , 300 , 1000 , + ∞ ) .
In these results, the biased-noise properties of these codes manifest themselves across the full landscape of possible pure and 2-Pauli biases, i.e. pure $X$, $Y$, and $Z$ noise, in addition to pure $XY$, $YZ$, and $XZ$ noise. In (a)-(c), we illustrate the bias tuning for pure $X$, $Y$, and $Z$ biases; consequently, as we tune the parameter $\eta$ in the direction towards zero, we effectively remove $X$, $Y$, and $Z$ noise from the simulations, in effect achieving, as $\eta\to 0$, pure $YZ$ noise from (a), pure $XZ$ noise from (b), and pure $XY$ noise from (c).
@@
Depolarizing $(1/3,1/3,1/3)$ $18.929$ $17.9\pm 0.81$ $18.98\pm 0.36$ $18.83\pm 0.13$ $18.46\pm 0.36$
Pure X $(1,0,0)$ $50.00$ $49.92\pm 0.096$ $10.79\pm 0.18$ $11.42\pm 0.52$ $33.56\pm 1.08$
Pure Y $(0,1,0)$ $50.00$ $49.91\pm 0.11$ $10.8\pm 0.19$ $44.66\pm 2.82$ $33.56\pm 1.08$
Pure Z $(0,0,1)$ $50.00$ $49.91\pm 0.11$ $10.8\pm 0.19$ $11.42\pm 0.52$ $22.99\pm 1.39$
Pure XZ $(1/2,0,1/2)$ $22.709$ $21.45\pm 0.68$ $22.1\pm 0.16$ $24.027\pm 1.08$ $22.08\pm 0.04$
Pure XY $(1/2,1/2,0)$ $22.709$ $21.45\pm 0.69$ $22.1\pm 0.17$ $21.77\pm 0.51$ $22.58\pm 0.17$
Pure YZ $(0,1/2,1/2)$ $22.709$ $21.45\pm 0.68$ $22.1\pm 0.17$ $21.77\pm 0.51$ $22.08\pm 0.04$
Table 2: Recovery threshold data points $p_{th}$ for select pure and 2-Pauli biases. The entry in blue surpasses the hashing bound; entries in green attain it up to statistical uncertainty, while entries in gold come within $2\%$ of achieving the bound.
@@
As is evidenced in each of the plots, the zero-rate HaPPY and tailored $\llbracket 7,1,3\rrbracket$ codes clearly achieve the hashing bound for all pure 1-Pauli biases, as well as for finite biases in which $\eta>10$. Additionally, in the limits for which $\eta\to 0$, the HaPPY and tailored $\llbracket 7,1,3\rrbracket$ code thresholds closely trace out the hashing bound as pure 2-Pauli noise is approached. The behavior of the HaPPY code emulates in large part the behavior seen of the generalized toric code family [24]; we discuss this in more detail in Section IV.
The Steane code portrays a distinct trend: in the pure 1-Pauli noise limit, the code performs significantly worse than all other codes tested. Nonetheless, the zero-rate Steane code eclipses the hashing bound for depolarizing noise ($\eta=0.5$), while also approaching the bound for finite biases as we approach the pure 2-Pauli limit ($\eta=0$). It was shown in previous work [23] that the color code exhibits similar behavior in the pure 1-Pauli bias regime; as such, it is logical to infer that the holographic Steane code manifests similar behavior in these limits. Due to the fact that both codes are built from self-dual CSS codes [42], in the 1-Pauli limit, only one set of stabilizers gives information about the error, while the other stabilizers do not yield additional information for decoding. As such, the problem reduces to a classical linear code with a check matrix defined by only half of the symplectic check matrix in the quantum code. Notwithstanding, it is known that many code concatenation schemes based on graph states can slightly exceed the hashing bound; our results, up to the uncertainty margin given, show that the holographic Steane can at least match most of the codes discovered in [27].
@@
In [21, 20], it was shown that randomized Clifford deformations, when applied to the surface code, can improve threshold performance with respect to a code-capacity noise channel. In this way, a surface code may be tailored to achieve the 1-Pauli hashing bound; one wonders whether such randomized approaches could be leveraged in holographic codes as well. Indeed, we may regard the additional Hadamard gates in the Evenbly code [56, 79] as an instance of Clifford deformation, in order to uphold the strict isometry properties present, while utilizing less highly-entangled quantum states.
As a preliminary step in this direction, we have performed a random Clifford-deformation of the holographic Steane code using Hadamard gates, with the goal of improving the 1-Pauli threshold for the code. The seed tensors used in this modified holographic code are Clifford-equivalent to that of the original Steane code; the only difference lies in the application of Hadamard gates to physical qubits $2,3,5$, and $6$. The resulting stabilizer generators and logical operators for this Clifford-deformed Steane code are given in Table 1.
As is shown above, the stabilizers and logical operators do not follow the typical convention as expected for mixing Pauli operators in a cyclic manner, nor as in generalized toric code constructions [24, 38, 25, 20]. The thresholds, reported in Figure 6, are for: (a) depolarizing noise ($19.0\pm 0.33\%$); (b) pure X noise ($\sim 50\%$); (c) pure Y noise ($10.58\pm 0.12\%$); and (d) pure Z noise ($\sim 50\%$). We observe that even our naive Clifford deformations, when combined with the holographic concatenation method, can greatly improve the threshold performance for targeting noise models of interest. Indeed, our results show that the Clifford-deformed holographic Steane code achieves the hashing bound for pure X and Z noise, and supersedes the bound for depolarizing noise. Such modifications will be investigated and are the subject of active research.
The SCF displays very interesting behavior at several points in Figure 5. For biased $X$ and $Z$ noise, the SCF performs similarly to the Steane code as $\eta\to+\infty$; this is due to the CSS properties of the seed tensor for both codes, as fewer parity checks are used for distinguishing errors in the limit of pure 1-Pauli noise. For biased $Y$ noise, however, the SCF exhibits excellent threshold performance; although the SCF does not attain the hashing bound in the pure $Y$ noise limit, many points of finite bias reach the bound. Additionally, as we tune the parameter $\eta$ back towards zero, the pure $XZ$ limit shows that the SCF surpasses the hashing bound at finite bias ($\eta\leq 0.1$). To the best of our knowledge, no other code construction thus far has surpassed the hashing bound for 2-Pauli noise [27, 15].

### v3

@@ II.3 Tensor-Network Decoding
Contracting a tensor network representing a holographic-code state with potentially thousands of qubits is computationally intractable; as such, we follow the simplification proposed in [farrelly_tn_codes, parallel_decoding_tn_codes]. In what follows, we briefly review this technique.
@@
As a side remark, the tensors of Equation 2 are not, strictly speaking, isometric. However, they facilitate efficient contraction, provided that the proper index contraction sequence is indicated; as such, the full decoding process which we will describe below exhibits a runtime complexity of $\mathcal{O}(n^{2.37})$ in the best case [bravyi2014efficient_tn_decoding, le2014powers]. More details on the formation of large tensor networks using the aforementioned scheme can be found in [farrelly_tn_codes, junyu_msc_thesis].
The tensor-network decoder is a maximum-likelihood (ML) decoder, which is shown to be optimal given the error model. The ML decoder generally works by calculating the most-likely correction needed to return the system to the correct code state, given a syndrome. Here, we define a syndrome as the set $\mathbf{s}=[s_{1}\cdots s_{n-k}]$ of $\mathbb{Z}_{2}$ parity results from measuring each stabilizer. An error $\mathsf{E}$ is initialized as a probability vector corresponding to each individual physical qubit; this vector is then contracted with the holographic tensor network, permitting us to calculate the syndrome $\mathbf{s}=\mathsf{HE}$, where $\mathsf{H}$ is the parity-check matrix of the holographic code.
@@
where $\text{Prob}[\mathsf{E}_{\text{pure}}|SL]$ signifies the probability of a pure error acting on the coset of stabilizer $S\in\mathsf{S}$ with respect to logical operator $L\in\mathsf{L}$. In practice, the decoder must find $L$ such that $\mathbb{P}(\mathsf{L},\mathbf{s})$ is maximized, i.e. $\operatorname*{arg\,max}\big[\mathbb{P}(\mathsf{L},\mathbf{s})\big]$. As an additional condition, the summation over all logical cosets must be equivalent to unity; that is $\sum_{L\in\mathsf{L}}\sum_{S\in\mathsf{S}}\mathbb{P}(\mathsf{L},\mathbf{s})=1$. As before, we refer the reader to [farrelly_tn_codes, parallel_decoding_tn_codes, junyu_msc_thesis] for further details on calculating the logical success rate.
@@ III.1 Noise Model and Setup
The error model tested in this work consists of the code-capacity error model for Pauli noise. Such an error model takes on the form
@@
[…]
@@
[…]
$\llbracket 6,1,3\rrbracket$ $ZIZIII\bar{I}$, $ZYYXI\bar{I}, XX$$XXXZI\bar{I}$, $IZZXIX\bar{I}$, $XYXYIZ\bar{I}$ $XZXZII\bar{X}$, $XYYXII\bar{Z}$
SCF $XXIXI\bar{I}$, $IIXXX\bar{I}$, $ZZI\bar{I}, IZIZ$$IZZ\bar{I}$ $XIXII\bar{X}$, $IIZIZ\bar{Z}$
CD-Steane $\bar{I}, XIZXZIII¯XIZXZI$$\bar{I}$, $XIIIZZX\bar{I}$, $\bar{I},ZIXZXIII¯ZIXZXI$$\bar{I}$, $ZIIIXXZ\bar{I}$ $ZZX\bar{X}, ZXXZXXZZ$$XXZ\bar{Z}$
@@
Threshold calculations were performed using the tensor-network decoder from [farrelly_tn_codes, parallel_decoding_tn_codes]. These decoders and others will be made public in an upcoming software package for holographic quantum error correction codes [fan2024lego_hqec]. However, for the purposes of understanding the current work, we focus on surveying biased-noise resilience for a swath of known holographic codes, and not on a systematic treatment of tensor-network decoding methodology [bravyi2014efficient_tn_decoding].
@@
In our biased-noise threshold profiling, we tested 43 distinct Pauli biases and have displayed them on the ternary diagram shown in Figure 2. Each point of the triangular plot shown represents a pure Pauli error channel. Subsequently, points on the interior of the plot can be read by following the grid lines provided to the boundaries; for example, the point to the lower left of the central depolarizing noise marker can be interpreted as having relative biases $\bar{r}=(r_{X},~r_{Y},~r_{Z})$ of (0.25,0.25,0.5)(0.25,~0.25,~0.5). At each bias point 10,000 Monte Carlo simulations were performed per threshold curve data point, per layer. We display commensurate threshold curve results for some select data points in Figure 4 for the zero-rate HaPPY code.
@@ III.2 Zero-Rate Holographic Code Results
Figure 3: Ternary plots for all holographic codes investigated in this study under biased noise, in the code-capacity setting. (a)-(d) depicts the zero-rate HaPPY, Steane, $\llbracket 6,1,3\rrbracket$, and SCF codes; thresholds are color-coded from dark blue ($p_{th}=10\%$) to dark red ($p_{th}=50\%$). Additionally, we tested the tailored $\llbracket 7,1,3\rrbracket$ code from [tailored_713, tailored_513_xzzx_codes], which achieved an identical threshold profile to that of the HaPPY code.
Figure 4: Threshold curves for depolarizing, pure $X$, $Y$, and $Z$ noise, as studied using the tensor-network decoder, for the zero-rate HaPPY code at up to $R=3$ layers of edge inflation.
@@
As a more detailed example, we plotted in Figure 4 the individual threshold curves obtained for the zero-rate HaPPY code for depolarizing noise, as well as for biased pure $X$, pure $Y$, and pure $Z$ noise. Around the fixed points in subfigures (b)-(d), the recovery rate promptly increases, which is due to several reasons: Firstly, the pure 1-Pauli noise capacity can be shown to be $50\%$. Secondly, as the tensor-network decoder is a maximum-likelihood decoder, after the $p=0.50$ mark, the decoder selects the most-probable error and returns a pure error (destabilizer) vector which is used to correct. Therefore, unlike a minimum weight decoder, the tensor network decoder is still able to provide the right correction with high probability even when $p>0.50$. More details can be found in [harris_iod, junyu_msc_thesis].
If we look beyond the pure 1-Pauli portion of the ternary plots, we note more subtle behavior of these codes: for example, it can be seen that modest threshold increases emerge for the Steane and $\llbracket 6,1,3\rrbracket$ codes as we move towards pure 2-Pauli noise, i.e., $XY$, YZYZ, and $XZ$ noise. These changes can be seen more clearly by examining the threshold behavior of the codes via the tuning of a bias parameter $\eta$. In previous work, this aim was accomplished by comparing various zero-rate codes against the hashing bound, a useful lower bound for quantifying code-capacity channel performance [xzzx_surface_code]. However, in contrast to the present work, all previous works have examined the pure 1-Pauli noise limit associated with the hashing bound, and have not delved into combinations of two types of Pauli errors at once, a condition known to occur in the square-lattice GKP code under a symmetric Gaussian random displacement noise channel [bosonic1, bosonic2, bosonic3, bosonic4], as well as for error channels resulting from performing Pauli twirling on the amplitude damping noise channel [amp_damping_twirl]. As such, we tune the $\eta$ parameter within the range $0\leq\eta\leq+\infty$, extending previous work past the regime indicated by $\eta\geq 0.5$ [xzzx_surface_code, clifford_deformed_surface, tailored_513_xzzx_codes].
@@
where $R$ represents an achievable rate $k/n$ for a random stabilizer code and $H(\bar{p})$ represents the Shannon entropy [bennett1998quantum]
[…]
@@
In these results, the biased-noise properties of these codes manifest themselves across the full landscape of possible pure and 2-Pauli biases, i.e. pure $X$, $Y$, and $Z$ noise, in addition to pure $XY$, YZYZ, and $XZ$ noise. In (a)-(c), we illustrate the bias tuning for pure $X$, $Y$, and $Z$ biases; consequently, as we tune the parameter $\eta$ in the direction towards zero, we effectively remove $X$, $Y$, and $Z$ noise from the simulations, in effect achieving, as $\eta\to 0$, pure $YZ$ noise from (a), pure $XZ$ noise from (b), and pure $XY$ noise from (c).
Noise bias $(r_{X},r_{Y},r_{Z})$ HaPPY ($\%$) Steane ($\%$) SCF ($\%$) $\llbracket 6,1,3\rrbracket$ ($\%$) Hashing ($\%$) TQEC ($\%$) Graph Codes ($\%$)
Depolarizing $(1/3,1/3,1/3)$ $17.9\pm 0.81$ $18.98\pm 0.36$ $18.83\pm 0.13$ $18.46\pm 0.36$ $18.929$ 19.14 $19.0597$
Pure X $(1,0,0)$ $49.92\pm 0.096$ $0.79\pm 0.18 11$$1.42\pm 0.52$ $33.56\pm 1.08$ $50.00$ 50.00 $50.00$
Pure Y $(0,1,0)$ $49.91\pm 0.11$ $10.8\pm 0.19$ $44.66\pm 2.82$ $33.56\pm 1.08$ $50.00$ $50.00$ $50.00$
Pure Z $(0,0,1)$ $49.91\pm 0.11$ $0.8\pm 0.19 11$$1.42\pm 0.52$ $22.99\pm 1.39$ $50.00$ $50.00$ $50.00$
Pure XZ $(1/2,0,1/2)$ $1.45\pm 0.68 22$$2.1\pm 0.16$ $4.027\pm 1.08 22$$08\pm 0.04 22.$$709 \approxHashing 22.$68422.684
Pure XY $(1/2,1/2,0)$ $1.45\pm 0.69 22$$2.1\pm 0.17$ $1.77\pm 0.51 22$$58\pm 0.17 22.$$709 \approxHashing 22.$68422.684
Pure YZ $(0,1/2,1/2)$ $1.45\pm 0.68 22$$2.1\pm 0.17$ $1.77\pm 0.51 22$$08\pm 0.04 22.$$709 \approxHashing 22.$68422.684
Table 2: Recovery threshold data points $p_{th}$ for select pure and 2-Pauli biases. The entry in blue surpasses the hashing bound; entries in green attain it up to statistical uncertainty, while entries in gold come within $2\%$ of achieving the bound.The “HaPPY” column refers to results obtained from both the original HaPPY code of [happy_paper], as well as those from concatenating the tailored $\llbracket 7,1,3\rrbracket$ [tailored_713]. In the last two columns, we compare to state-of-the-art results for topological codes (TQEC) as well as for graph codes: Depolarizing-noise thresholds for hexagonal color codes are taken from [bombin_strong] and pure Pauli noise in surface codes from [tuckett_biased_surface, xzzx_surface_code], where thresholds around the Hashing bound were observed near the two-Pauli noise regime (the exact zero-bias regimes were not considered); graph code results are for the 5-in-5 code from [leditzky_pauli_hashing].
@@
As is evidenced in each of the plots, the zero-rate HaPPY and tailored $\llbracket 7,1,3\rrbracket$ codes clearly achieve the hashing bound for all pure 1-Pauli biases, as well as for finite biases in which $\eta>10$. Additionally, in the limits for which $\eta\to 0$, the HaPPY and tailored $\llbracket 7,1,3\rrbracket$ code thresholds closely trace out the hashing bound as pure 2-Pauli noise is approached. The behavior of the HaPPY code emulates in large part the behavior seen of the generalized toric code family [tailored_513_xzzx_codes]; we discuss this in more detail in Section IV.
The Steane code portrays a distinct trend: in the pure 1-Pauli noise limit, the code performs significantly worse than all other codes tested. Nonetheless, the zero-rate Steane code eclipses the hashing bound for depolarizing noise ($\eta=0.5$), while also approaching the bound for finite biases as we approach the pure 2-Pauli limit ($\eta=0$). It was shown in previous work [tuckett_biased_surface] that the color code exhibits similar behavior in the pure 1-Pauli bias regime; as such, it is logical to infer that the holographic Steane code manifests similar behavior in these limits. Due to the fact that both codes are built from self-dual CSS codes [lidar_qec], in the 1-Pauli limit, only one set of stabilizers gives information about the error, while the other stabilizers do not yield additional information for decoding. As such, the problem reduces to a classical linear code with a check matrix defined by only half of the symplectic check matrix in the quantum code. Notwithstanding, it is known that many code concatenation schemes based on graph states can slightly exceed the hashing bound; our results, up to the uncertainty margin given, show that the holographic Steane can at least match most of the codes discovered in [leditzky_pauli_hashing].
@@
In [3d_surface_hashing, clifford_deformed_surface], it was shown that randomized Clifford deformations, when applied to the surface code, can improve threshold performance with respect to a code-capacity noise channel. In this way, a surface code may be tailored to achieve the 1-Pauli hashing bound; one wonders whether such randomized approaches could be leveraged in holographic codes as well. Indeed, we may regard the additional Hadamard gates in the Evenbly code [evenbly_codes2, steinberg_evenbly_codes1] as an instance of Clifford deformation, in order to uphold the strict isometry properties present, while utilizing less highly-entangled quantum states.
As a preliminary step in this direction, we have performed a random Clifford deformation of the holographic Steane code using Hadamard gates, with the goal of improving the 1-Pauli threshold for the code. The seed tensors used in this modified holographic code are Clifford-equivalent to that of the original Steane code; the only difference lies in the application of Hadamard gates to physical qubits $2,3,5$, and $6$. The resulting stabilizer generators and logical operators for this Clifford-deformed Steane code are given in Table 1.
As is shown above, the stabilizers and logical operators do not follow the typical convention as expected for mixing Pauli operators in a cyclic manner, nor as in generalized toric code constructions [tailored_513_xzzx_codes, tailored_713, xzzx_surface_code, clifford_deformed_surface]. The thresholds, reported in Figure 6, are for: (a) depolarizing noise ($19.0\pm 0.33\%$); (b) pure X noise ($\sim 50\%$); (c) pure Y noise ($10.58\pm 0.12\%$); and (d) pure Z noise ($\sim 50\%$). We observe that even our naive Clifford deformations, when combined with the holographic concatenation method, can greatly improve the threshold performance for targeting noise models of interest. Indeed, our results show that the Clifford-deformed holographic Steane code achieves the hashing bound for pure X and Z noise, and supersedes the bound for depolarizing noise. Such modifications will be investigated and are the subject of active research.
The SCF displays very interesting behavior at several points in Figure 5. For biased $X$ and $Z$ noise, the SCF performs similarly to the Steane code as $\eta\to+\infty$; this is due to the CSS properties of the seed tensor for both codes, as fewer parity checks are used for distinguishing errors in the limit of pure 1-Pauli noise. For biased $Y$ noise, however, the SCF exhibits excellent threshold performance; although the SCF does not attain the hashing bound in the pure $Y$ noise limit, many points of finite bias reach the bound. Additionally, as we tune the parameter $\eta$ back towards zero, the pure $XZ$ limit shows that the SCF surpasses the hashing bound at finite bias ($\eta\leq 0.1$). To the best of our knowledge, no other code construction thus far has surpassed the hashing bound for 2-Pauli noise [leditzky_pauli_hashing, lower_bounds_capacity_birgitta].